lim [(x-3)/(x-1)]^x = 1/e^2


x→∞

1
Expert's answer
2022-01-10T11:22:37-0500
"\lim\limits_{x\to\infin}\bigg(\dfrac{x-3}{x-1}\bigg)^x=\lim\limits_{x\to\infin}\bigg(1-\dfrac{2}{x-1}\bigg)^x"

"=\lim\limits_{x\to\infin}\bigg(1-\dfrac{2}{x-1}\bigg)^{-{x-1 \over 2}(-{2 \over x-1})x}"

"=\lim\limits_{x\to\infin}\bigg[\bigg(1-\dfrac{2}{x-1}\bigg)^{-{x-1 \over 2}}\bigg]^{-{2x \over x-1}}"

"=e^{-2}=\dfrac{1}{e^2}"